OpenStudy (anonymous):

If an 10 kg object has a potential energy of 125 J, what velocity would it be right before it hit the ground?

OpenStudy (poofypenguin):

By potential energy do you mean gravitational potential energy, as in the object is falling?

OpenStudy (anonymous):

i think it is while the object is falling

OpenStudy (poofypenguin):

The gravitational potential energy is being transferred into kinetic energy so, Eg=Ek Eg=0.5m(v^2) 125 J = 0.5*(10 kg)*(v^2) 125 J = 5 kg * (v^2) 125J/ 5kg = v^2 25=v^2 5 m/s = v